# Strong and Noise-Tolerant Entanglement in Dissipative Optomechanics

### v1

@@ Abstract
Macroscopic entanglement, as critical quantum resources in quantum information science, has been extensively studied in optomechanical systems with purely dispersive coupling over the past decades. However, quantum entanglement, induced by purely dissipative coupling, remains unexplored. In this work, we study quantum entanglement in a Michelson–Sagnac interferometer, where the dispersive and the dissipative coupling can be arbitrarily switched on and off. With current experimental parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power with the purely dispersive (dissipative) coupling. Further, we find that quantum entanglement generated by purely dissipative coupling is significantly stronger and more noise-tolerant than that generated by purely dispersive coupling. When both couplings coexist, entanglement is weakened due to interference. Our findings provide a promising path to engineer strong and noise-tolerant quantum entanglement in purely dissipative quantum systems.
@@ I Introduction
Due to its potential applications in gravitational wave detection Abbott et al. (2016) and quantum-classical boundary tests Haroche (1998), cavity optomechanics has garnered significant theoretical and experimental interest in quantum information science Aspelmeyer et al. (2014). With advances in ground-state cooling Liu et al. (2013); Teufel et al. (2011); Gröblacher et al. (2009); Schliesser et al. (2008), exploring quantum effects in cavity optomechanics has become a reality Lü et al. (2015); Safavi-Naeini et al. (2013); Purdy et al. (2013); Weis et al. (2010); Safavi-Naeini et al. (2011); Kronwald and Marquardt (2013); Koppenhöfer et al. (2023); Guo et al. (2016); Kyriienko et al. (2014a); Xiong et al. (2016, 2021a); Chen et al. (2021); Lü et al. (2013); Peng et al. (2023); Xiong et al. (2023); Wang et al. (2024); Mann et al. (2018); Jäger et al. (2019); Xiong et al. (2021b, 2022), especially for macroscopic quantum entanglement Vitali et al. (2007). Over the past two decades, numerous theoretical proposals have been put forward to achieve optomechanical entanglement Tian (2012, 2013); Wang and Clerk (2013); Chen et al. (2024, 2023); Liu et al. (2024a); Lai et al. (2022); Jiao et al. (2020, 2022); Liu et al. (2024b). Experimentally, macroscopic entanglement between two massive oscillators has also been realized in cavity optomechanics experimentally Ockeloen-Korppi et al. (2018). Notably, both theoretical and experimental proposals have predominantly relied on coherent dispersive coupling mechanisms between the cavity mode and the mechanical mode.
@@
This paper is organized as follows. In Sec. II, we introduce the proposed model and its Hamiltonian. Then, the steady-state solutions of the system are investigated in Sec. III. In Sec. IV, the dynamics of the fluctuation and covariance matrix are given. In Secs. V, quantum entanglement is investigated. Finally, a conclusion is given in Sec. VI.
@@ II The model and Hamiltonian
where $\omega_{m}$ is the frequency of the membrane, $\omega$ is the frequency of the bosonic mode (i.e., bath) with annihilation (creation) operator $a_{\omega}$ ($a_{\omega}^{\dagger}$), $x^{\prime}=\sqrt{2\kappa_{a}}-{g_{\kappa}x}/{\sqrt{2\kappa_{a}}}$ is the position-dependent coupling strength between the optical cavity and the bosonic mode. In fact, the bosonic mode here is to introduce a zero-value expectation input field $a_{\rm in}=\int\frac{d\omega}{\sqrt{2\pi}}e^{-i\omega(t-t_{0})}a_{\omega,0}$, with $a_{\omega,0}$ the operator at an initial time, to the cavity mode $a$, which can be easily obtained following the standard input-output theory with the Markov approximation. At the optical domain, the input field satisfies the relations walls1994gj
@@ VI Conclusion
In this paper, we investigate quantum entanglement in a dissipative optomechanical system, acted by a fixed mirror and a Michelson–Sagnac interferometer with a membrane. This system allows the dissipative and dispersive couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the beam splitter, enabling us to explore quantum entanglement in three distinct scenarios: purely dispersive case, purely dissipative case, and combined coupling case. Our analysis reveals that, in the purely dispersive case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, whereas in the purely dissipative case, the displacement shows a linear dependence. Furthermore, entanglement generated through purely dissipative coupling is not only significantly stronger but also more noise-tolerant compared to that achieved via purely dispersive coupling. When both coupling mechanisms coexist, quantum entanglement is reduced, likely due to the interference between the two coupling interactions. These findings suggest the potential of purely dissipative coupling to engineer strong and noise-tolerant quantum entanglement, offering a promising pathway for advancing quantum technologies.
This work is supported by the National Natural Science Foundation of China (Grant No. 12175001 ), the “Pioneer” and “Leading Goose” R&\&&D Program of Zhejiang (Grant No. 2025C01028), the Natural Science Foundation of Zhejiang Province (Grant No. LY24A040004), and the Shenzhen International Quantum Academy (Grant No. SIQA202412).

### v2

@@ Abstract
Macroscopic entanglement, as critical quantum resources in quantum information science, has been extensively studied in optomechanical systems with purely dispersive coupling over the past decades. However, quantum entanglement, induced by purely dissipative coupling, remains unexplored. In this work, we study quantum entanglement in a Michelson–Sagnac interferometer, where the dipersive and the dissipative coupling can be arbitrarily switched on and off. With current experimental parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power with the purely dispersive (dissipative) coupling. Further, we find that quantum entanglement generated by purely dissipative coupling is significantly stronger and more noise-tolerant than that generated by purely dispersive coupling. When both couplings coexist, entanglement is weakened due to interference. Our findings provide a promising path to engineer strong and noise-tolerant quantum entanglement in purely dissipative quantum systems.
@@ I Introduction
Due to its potential applications in gravitational wave detection Abbott et al. (2016) and quantum-classical boundary tests Haroche (1998), cavity optomechanics has garnered significant theoretical and experimental interest in quantum information science Aspelmeyer et al. (2014). With advances in ground-state cooling Liu et al. (2013); Teufel et al. (2011); Gröblacher et al. (2009); Schliesser et al. (2008), exploring quantum effects in cavity optomechanics has become a reality Lü et al. (2015); Safavi-Naeini et al. (2013); Purdy et al. (2013); Weis et al. (2010); Safavi-Naeini et al. (2011); Kronwald and Marquardt (2013); Koppenhöfer et al. (2023); Guo et al. (2016); Kyriienko et al. (2014a); Xiong et al. (2016, 2021a); Chen et al. (2021); Lü et al. (2013); Peng et al. (2023); Xiong et al. (2023); Wang et al. (2024); Mann et al. (2018); Jäger et al. (2019); Xiong et al. (2021b, 2022), especially for macroscopic quantum entanglement Vitali et al. (2007). Over the past two decades, numerous theoretical proposals have been put forward to achieve optomechanical entanglement Tian (2012, 2013); Wang and Clerk (2013); Chen et al. (2024, 2023); Liu et al. (2024a); Lai et al. (2022); Jiao et al. (2020, 2022); Liu et al. (2024b); Shang and Li (2024). Experimentally, macroscopic entanglement between two massive oscillators has also been realized in cavity optomechanics experimentally Ockeloen-Korppi et al. (2018). Notably, both theoretical and experimental proposals have predominantly relied on coherent dispersive coupling mechanisms between the cavity mode and the mechanical mode.
@@
The rest paper is organized as follows. In Sec. II, we introduce the proposed model and its Hamiltonian. Then, the steady-state solutions of the system are investigated in Sec. III. In Sec. IV, the dynamics of the fluctuation and covariance matrix are given. In Secs. V, quantum entanglement is investigated. Finally, a conclusion is given in Sec. VI.
@@ II The model and Hamiltonian
where $\omega_{m}$ is the frequency of the membrane, $\omega$ is the frequency of the bosonic mode (i.e., bath) with annihilation (creation) operator $a_{\omega}$ ($a_{\omega}^{\dagger}$), $x^{\prime}=\sqrt{2\kappa_{a}}-{g_{\kappa}x}/{\sqrt{2\kappa_{a}}}$ is the position-dependent coupling strength between the optical cavity and the bosonic mode. In fact, the bosonic mode here is to introduce a zero-value expectation input field $a_{\rm in}=\int\frac{d\omega}{\sqrt{2\pi}}e^{-i\omega(t-t_{0})}a_{\omega,0}$, with $a_{\omega,0}$ the operator at an initial time, to the cavity mode $a$, which can be easily obtained following the standard input-output theory with the Markov approximation. At the optical domain, the input field satisfies the relations Walls and Milburn (1994)
@@ VI Conclusion
In this paper, we investigate quantum entanglement in a dissipative optomechanical system, acted by a fixed mirror and a Michelson–Sagnac interferometer with a membrane. This system allows the dissipative and dispersive couplings to be independently tuned on and off by adjusting the membrane position or the reflectivity (transmissivity) of the BS, enabling us to explore quantum entanglement in three distinct scenarios: purely dispersive case, purely dissipative case, and combined coupling case. Our analysis reveals that, in the purely dispersive case, the steady-state mechanical displacement exhibits a nonlinear dependence on the driving power, whereas in the purely dissipative case, the displacement shows a linear dependence. Furthermore, entanglement generated through purely dissipative coupling is not only significantly stronger but also more noise-tolerant compared to that achieved via purely dispersive coupling. When both coupling mechanisms coexist, quantum entanglement is reduced, likely due to the interference between the two coupling interactions. These findings suggest the potential of purely dissipative coupling to engineer strong and noise-tolerant quantum entanglement, offering a promising pathway for advancing quantum technologies.
This work is supported by the National Natural Science Foundation of China (Grant No. 12175001 ), the “Pioneer” and “Leading Goose” R&\&&D Program of Zhejiang (Grant No. 2025C01028), the Natural Science Foundation of Zhejiang Province (Grant No. LY24A040004), and the Shenzhen International Quantum Academy (Grant No. SIQA2024KFKT010).